### v1

@@ Abstract
Since the advent of the “Neural Ordinary Differential Equation (Neural ODE)” paper[1], learning ODEs with deep learning has been applied to system identification, time-series forecasting, and related areas. Exploiting the diffeomorphic nature of ODE solution maps, neural ODEs has also enabled their use in generative modeling. Despite the rich potential to incorporate various kinds of physical information, training Neural ODEs remains challenging in practice. This study demonstrates, through the simplest one-dimensional linear model, why training Neural ODEs is difficult. We then propose a new stabilization method and provide an analytical convergence analysis. The insights and techniques presented here serve as a concise tutorial for researchers beginning work on Neural ODEs.
@@
Neural ODE, Gradient Method, Optimization
@@ 1 Introduction
$(x_{t},y_{t},z_{t})\leftarrow\left(x_{t},\tfrac{\partial{\mathcal{L}}}{\partial x_{t}},{\bf 0}\right)$triangleright Initialize augmented state
$(x_{0},y_{0},z_{0})\leftarrow\texttt{ODESolver}\bigg{(}\Big{(}-f_{\theta},-\tfrac{\partial f_{\theta}}{\partial x}^{\mathsf{T}}y,-y^{\mathsf{T}}\tfrac{\partial f_{\theta}}{\partial\theta}\Big{)},(x_{t},y_{t},z_{t}),t\bigg{)}$triangleright Solve Reversal-time ODE
[…]
return $z_{0}$triangleright Return gradient
@@ 3 Background
where $p$ is the distribution of initial conditions.
@@ 4 Problem
     ${\mathcal{G}},{\mathcal{X}}_{t}\leftarrow\{\},\{\}$triangleright Initialize gradient and terminal state set
@@
         ${\mathcal{X}}_{t}.\text{append}(x_{t})\quad{\mathcal{G}}.\text{append}(g)$triangleright Save each terminal state and gradient set
@@
     $\sigma^{2}_{t}\leftarrow\tfrac{1}{N}\sum_{x_{t}\in{\mathcal{X}}_{t}}x_{t}^{2}$triangleright Calculate sample terminal variance
     $\nabla{\rm Loss}\leftarrow\tfrac{1}{N}\sum_{g\in{\mathcal{G}}}g$triangleright Calculate sample mean gradient
     $\theta\leftarrow\theta-\eta\tfrac{1}{t^{2}\sigma^{2}_{t}}\nabla{\rm Loss}$triangleright Update parameters
@@
$\hskip 20.00003ptf_{\theta}(x)\triangleq ax,\quad f_{*}(x)\triangleq a^{*}x,$ (4)
and assume the initial distribution $p$ is Gaussian with mean 00 and variance $\sigma^{2}$. Then the loss (2) becomes
@@ 5 Method
(a) Loss (SGD)
(b) Loss (Adam)
(c) Loss (Our method)
(d) Parameter (SGD)
(e) Parameter (Adam)
(f) Parameter (Our method)
@@
[…]
@@ 6 Experiments
To verify effectiveness, we trained the one-dimensional linear dynamics $f_{\theta}(x)=ax$ toward the target $f_{*}(x)=-x$ with learning rate $\eta=0.5$. Baselines were stochastic gradient descent (SGD) and Adam [12] with the same learning rate.
@@ 7 Discussion
Non-uniqueness of dynamics for a given endpoint map: Even with Lipschitz continuity, the dynamics $f$ are not uniquely determined by the solution map $\phi^{t}_{f}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$. Linear multi-dimensional systems possess rotational degrees of freedom; nonlinear systems admit continuous freedom in velocity fields, complicating optimization. Uniqueness requires additional conditions, such as temporal sampling or structural constraints.
Difficulty of per-parameter gradient scaling: Nonlinear dynamics involve many parameters, making gradient scaling challenging. While methods like natural gradient descent [13] account for parameter influence, Neural ODEs are further complicated by the intervening solver. The present variance-based scaling applies neatly only to commutative, symmetric linear dynamics.
@@ 8 Conclusion
Using a one-dimensional linear system, we (1) elucidated the source of training difficulty in Neural ODEs, (2) proposed a variance-corrected gradient rule with proven linear convergence, and (3) validated the theory on a toy example. Although limited to a trivial setting, the insights serve as a stepping stone for future Neural ODE research.

### v3

@@ Abstract
Since the advent of the “Neural Ordinary Differential Equation (Neural ODE)” paper[1], learning ODEs with deep learning has been applied to system identification, time-series forecasting, and related areas. Exploiting the diffeomorphic nature of ODE solution maps, neural ODEs have also enabled their use in generative modeling. Despite the rich potential to incorporate various kinds of physical information, training Neural ODEs remains challenging in practice. This study demonstrates, through the simplest one-dimensional linear model, why training Neural ODEs is difficult. We then propose a new stabilization method and provide an analytical convergence analysis. The insights and techniques presented here serve as a concise tutorial for researchers beginning work on Neural ODEs.
@@
Neural ODE, Gradient Method, Non-Convex Optimization
@@ 1 Introduction
$(x_{t},y_{t},z_{t})\leftarrow\left(x_{t},\tfrac{\partial{\mathcal{L}}}{\partial x_{t}},{\bf 0}\right)$ $\triangleright$ Initialize augmented state
$(x_{0},y_{0},z_{0})\leftarrow\texttt{ODESolver}\bigg{(}\Big{(}-f_{\theta},-\tfrac{\partial f_{\theta}}{\partial x}^{\mathsf{T}}y,-y^{\mathsf{T}}\tfrac{\partial f_{\theta}}{\partial\theta}\Big{)},(x_{t},y_{t},z_{t}),t\bigg{)}$ $\triangleright$ Solve Reversal-time ODE
/* $\texttt{ODESolver}(\text{dynamics},\text{initial stete},\text{time})$ is an numerical ODE integrator */
return $z_{0}$ $\triangleright$ Return gradient
@@ 3 Background
where $p$ is the distribution of initial state.
@@ 4 Problem
  ${\mathcal{G}},{\mathcal{X}}_{t}\leftarrow\{\},\{\}$ $\triangleright$ Initialize gradient and terminal state set
@@
   ${\mathcal{X}}_{t}.\text{append}(x_{t})\quad{\mathcal{G}}.\text{append}(g)$ $\triangleright$ Save each terminal state and gradient
@@
  $\sigma^{2}_{t}\leftarrow\tfrac{1}{N}\sum_{x_{t}\in{\mathcal{X}}_{t}}x_{t}^{2}$ $\triangleright$ Calculate sample terminal variance
  $\nabla{\rm Loss}\leftarrow\tfrac{1}{N}\sum_{g\in{\mathcal{G}}}g$ $\triangleright$ Calculate sample mean gradient
  $\theta\leftarrow\theta-\eta\tfrac{1}{t^{2}\sigma^{2}_{t}}\nabla{\rm Loss}$ $\triangleright$ Update parameters
@@
$\hskip 20.00003ptf_{\theta}(x)\triangleq ax,\quad f_{*}(x)\triangleq a^{*}x$ (4)
and assume the initial distribution $p$ is Gaussian with mean 0 and variance $\sigma^{2}$. Then the loss (2) becomes
@@ 5 Method
(a) SGD
(b) Adam
(c) Natural Gradient (Theorem 1)
@@
Whereas standard gradient descent depends on $t$ and $\sigma_{k}^{2}(t)$, our correction removes this dependence, guaranteeing linear convergence of the loss to zero. Algorithm 2 provides the pseudocode for the aforementioned gradient method. One can see that $\sigma_{k}^{2}(t)$ is obtained naturally when deriving the gradient with the adjoint method.
@@ 6 Experiments
To verify effectiveness, we trained the one-dimensional linear dynamics $f_{\theta}(x)=ax$ toward the target $f_{*}(x)=-x$ with learning rate $\eta=0.05$. Baselines were stochastic gradient descent (SGD) and Adam [12] with the same learning rate.
@@ 7 Discussion
Theorem 5.1 as the Natural Gradient Method:
So, what was the correction term proposed in our Theorem? This actually corresponds to the inverse of the Fisher Information Matrix:
$F(a)$ $\triangleq{\mathrm{E}}\left[\left(\nabla\phi_{f_{\theta}}^{t}(x)\right)\left(\nabla\phi_{f_{\theta}}^{t}(x)\right)^{\mathsf{T}}\right]$
$={\mathrm{E}}\left[(te^{at}x)^{2}\right]=t^{2}\sigma^{2}_{k}(t)$
Therefore, Theorem 5.1 reduces to nothing more than the natural‑gradient method [13] for the simplest ODE. This suggests that training a Neural ODE appears to be a problem on which the only natural‑gradient method can be exceptionally efficient.
On the other hand, applying the natural gradient method to the dynamics $f_{\theta}$ of a general neural network is difficult. The computation of the inverse of the Fisher Information Matrix requires a computational complexity of at least $O(n^{2})$, where $n$ is the number of parameters. Consequently, efficiently training neural networks for deep learning, which involves a vast number of parameters, remains a challenge.
Non-uniqueness of dynamics for a given endpoint map:
Even with Lipschitz continuity, the dynamics $f$ are not uniquely determined by the solution map $\phi^{t}_{f}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$. Linear multi-dimensional systems possess rotational degrees of freedom; nonlinear systems admit continuous freedom in velocity fields, complicating optimization. Uniqueness requires additional conditions, such as temporal sampling or structural constraints.
@@ 8 Conclusion
Using a one-dimensional linear system, we (1) elucidated the source of training difficulty in Neural ODEs, (2) proposed a variance-corrected gradient rule with proven linear convergence, and (3) validated the theory on a toy example. Furthermore, we confirm that this gradient-based method is equivalent to the natural gradient method for Neural ODEs. Although limited to a trivial setting, the insights serve as a stepping stone for future Neural ODE research.